# Orthogonal projections of hypercubes

### v3

@@ Abstract
Projections of hypercubes have been applied to visualize high-dimensional binary state spaces in various scientific fields. Conventional methods for projecting hypercubes, however, face practical difficulties. Manual methods require nontrivial adjustments of the projection basis, while optimization-based algorithms limit the interpretability and reproducibility of the resulting plots. These limitations motivate us to explore theoretically analyzable projection algorithms such as principal component analysis (PCA). Here, we investigate the mathematical properties of PCA-projected hypercubes. Our numerical and analytical results show that PCA effectively captures polarized distributions within the hypercubic state space. This property enables the assessment of the asymptotic distribution of projected vertices and error bounds, which characterize the performance of PCA in the projected space. We demonstrate the application of PCA to visualize the hypercubic energy landscapes of Ising spin systems. By adding projected hypercubic edges, these visualizations reveal pathways of correlated spin flips. Our work provides a better understanding of how PCA discovers hidden patterns in high-dimensional binary data.
@@ I Introduction
Human interest in higher dimensions, particularly the fourth dimension, has been emerging since the 19th century [1, 2, 3, 4]—before the theory of relativity by A. Einstein [5, 6, 7], which is well known as the theory of four-dimensional spacetime [8]. Unlike the theory of relativity, 19th-century mathematicians, physicists, and philosophers focused on the spatial fourth dimension. Among them, C. H. Hinton has been influential because his works [9] indicate that the tesseract, or four-dimensional cube, can embody four-dimensional space. Since then high-dimensional space has been explored with hypercubes.
@@
By generalizing the idea of a three-dimensional cube to higher dimensions, one obtains a high-dimensional cube or a hypercube [10] (see Appendix A for details on the construction of a hypercube). Although a hypercube is purely a geometric concept [11, 12, 10, 13], it has applications not only in physics [14, 15, 16], electrical engineering [17, 18, 19, 20, 21, 22], graph theory [23, 18], and recreational mathematics [18, 13, 24], but also in a wide range of interdisciplinary fields [25]. In various arts, e.g., literature [26, 27, 24], visual art [28, 1, 29, 30], architecture [28], ornament [31], and film [32, 33, 34, 35], hypercubes have repeatedly inspired human imagination and have become an embodiment of high-dimensional space. Since the pioneering work of almost 150 years ago [11] and subsequent contributions [9, 10], interest in high-dimensional geometry, especially hypercubes, has been continuously growing.
@@
Returning to the sciences, hypercubes visualize binary state111 Also called microstate, configuration, or phase. space in diverse fields, ranging from physics to biology.222 See, for instance, evolutionary landscapes [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], epistasis [36, 48, 49], chemical reaction networks [50, 51], learning in neural networks [52], genetic code space [53, 54, 45], allostery [55, 56, 57, 58, 59, 60], quantum states [61, 62], quantum walks [63], language space [64], data visualization [65, 66], probability currents [67], protein folding [68, 69], energy landscapes [70], gene regulation [71, 72, 73, 74, 75], quantum many-body scar states [76, 77, 78], morphology [79, 80], disease progression [45, 46], gene regulatory networks [81], gene expression [82], and spike dynamics [83, 84]. This is because hypercubic vertices correspond to binary states of a system [Figs. 1(a) and 1(b)], and hypercubic edges represent transitions333 Strictly speaking, only one of the dimensions, components, or elements of the system must change in the transition on the hypercubic edge. This kind of dynamics is called asynchronous dynamics (update) [85, 86, 87, 81]. It is also called multipartite dynamics [88, 89, 83, 84] in the context of stochastic thermodynamics. On the other hand, transitions with several changes of elements are called synchronous dynamics (update) [90, 91, 92, 93, 94, 95, 96]. between states [Figs. 1(c) and 1(d)]. This can be naturally applied to illustrate the high-dimensional binary state space or state transition diagram. We refer to such a binary state space as a hypercubic state space. As an example of transitions between the vertices, we visualize the Hamiltonian path on a hypercube [Figs. 1(c) and 1(d), Tables 1 and 2].
@@
An alternative method to project a hypercube is optimizing the coordinates of projected vertices following a predefined error function [97]. While this method projects hypercubes automatically, it has several limitations. The first limitation is reproducibility. Because of the stochastic nature of the optimization process, projections generated from the same vertices are not guaranteed to be the same (unless all relevant parameters are fixed). Interpretability is another weakness of projecting hypercubes by optimization. Because of its nonlinear nature, what is indicated from the resulting plot is not obvious (see, however, [98, 99, 100]). How can one create reproducible and interpretable projections of hypercubes?
Here, we reveal that principal component analysis (PCA) [101, 102] provides reproducible, interpretable, and automatic projections of hypercubes. PCA is a linear dimensionality reduction method frequently performed in statistics and machine learning [103, 104]. By interpreting the principal component (PC) loading as the basis for the projection of a hypercube, we show that one can draw the edges of a hypercube on a projected two-dimensional plane. This idea is closely related to the biplot [105], a method to visualize loadings (eigenvectors) with data points to assist in the interpretation of resulting plots. By combining ideas from geometry and statistics (machine learning), we achieve informative projections of hypercubes [79]: we can analytically obtain the properties of the projection, and the resulting plots are examined with such prior knowledge.
This paper is organized as follows. In Sec. II, we present the Hamming and fractal projections with introducing manual orthogonal projections of hypercubes. Then, in Sec. III, we show how biplots enable us to interpret the resulting plots of PCA on hypercubic vertices. We provide some examples of orthogonal projections of hypercubes using PCA in Sec. IV. Through analytical and numerical investigations in Sec. V, we show that principal components are informative for weighted vertices. In Sec. VI, using the inner-product error, we show that the vertices around the origin of the projected space less accurately preserve the original distances between them. We apply our method to visualize the hypercubic energy landscape [70] of Ising spin systems in Sec. VII, and conclude this paper in Sec. VIII.
@@ II Orthogonal Projections of Hypercubes
Orthogonal projection is a linear method to project a high-dimensional object. We introduce the concept of the contribution basis, which is the projected unit vector of each dimension. We then illustrate isometric, Hamming, and fractal projections of hypercubes by manually varying the contribution basis.
@@ II.1 Dimensionality reduction and orthogonal projection
where $\bm{b}\coloneqq\begin{bmatrix}b_{1}&\cdots&b_{N}\end{bmatrix}^{\top}\in\left\{1,0\right\}^{N}$ is an $N$-dimensional coordinate of the hypercubic vertices. The reduced original dimensions are represented by two $N$-dimensional vectors, $\bm{v}_{1},\bm{v}_{2}\in\mathbb{R}^{N}$.
II.2 Contribution basis
@@
$\bm{r}=\sum_{i=1}^{N}b_{i}\bm{e}_{i},$ (2)
where the contribution basis,
$\bm{e}_{i}\coloneqq\begin{bmatrix}v_{i;1}\\ v_{i;2}\end{bmatrix},$ (3)
@@
[$\ \bm{e}_{1}&\cdots&\bm{e}_{N}\\ \vrule&&\vrule\end{bmatrix}\coloneqq\begin{bmatrix}\text{---}&\bm{v}_{1}^{\top}&\text{---}\\ \text{---}&\bm{v}_{2}^{\top}&\text{---}\end{bmatrix}.$ (4)
@@
The interpretation of the contribution basis in Eq. (2) is the contribution from the $i$th dimension to the projected space. Because $b_{i}\in\left\{1,0\right\}$, the contribution basis corresponds to the projected unit vector of each dimension. Thus, by introducing the contribution basis, we can decompose the resulting projection into the contributions from each dimension as shown in Eq. (2). The remaining question is how to determine the contribution basis $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$.
Figure 2: A gallery of orthogonal projections of hypercubes. Colored arrows represent the contribution basis corresponding to the original dimensions. The boxes on the bottom right indicate the correspondence between the colors and the original dimensions. (a) An isometric projection of a cube. Notice that $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$ are overlapped. (b) An isometric projection of a tesseract. (c) A Hamming projection of a cube. (d) A Hamming projection of a tesseract. Notice that $\begin{bmatrix}1&0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1&0\end{bmatrix}^{\top}$ are overlapped. In Hamming projections (c) and (d), the contribution basis of each dimension has the same horizontal contribution. (e) A fractal projection of a six-dimensional hypercube. (f) A fractal projection of an eight-dimensional hypercube. In fractal projections (e) and (f), the contribution basis of the first half (left half) of the code are ten times longer than the rest. More projections of hypercubes are available in the Supplemental Material [109].
@@ II.4 Hamming and fractal projections
By modifying the contribution basis $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$ for projection, one can view a high-dimensional hypercube from various angles. Several methods reflecting the cross-section of the hypercube have already been suggested [10]. Here, we introduce two other orthogonal projections of hypercubes used in sciences.
A Hamming projection [Figs. 2(c) and 2(d)], we call, is a type of projection that plots vertices according to the Hamming distance [110] from a selected vertex. To achieve this, an example of a contribution basis is $\bm{e}_{i}=\begin{bmatrix}1&\left(i-\frac{N+1}{2}\right)\frac{2}{N-1}\sin\left(\frac{\varphi}{2}\right)\end{bmatrix}^{\top}$, where $\varphi\coloneqq\arccos\left(\frac{\bm{e}_{1}^{\top}\bm{e}_{N}}{\sqrt{\bm{e}_{1}^{\top}\bm{e}_{1}}\sqrt{\bm{e}_{N}^{\top}\bm{e}_{N}}}\right)$ is the angle between the first and last contribution basis, but it can be any. The first element of $\bm{e}_{i}$ is the same for all $i$. The second element can be any, but we determine them to increase linearly with $i$. This type of projection is found not only in research in physics444 Examples are found in quantum walk [63], unsupervised learning of states of Ising spin system [111, 112, 113], and quantum many-body scar states [76, 77, 78]. but also in biology.555 Examples include evolutionary landscapes [36, 40, 41, 42, 43, 44, 46, 47], epistasis [36, 48], genetic code space [45], data visualization [65, 66], protein folding [68, 69], gene regulation [71], disease progression [45, 46], and allostery [59]. The horizontal distance between the Hamming projected hypercubic vertices corresponds to the Hamming distance from a reference vertex on the rightmost or leftmost side of the projected space. In Figs. 2(c) and 2(d), the horizontal axis corresponds to the Hamming distance from the origin $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&0&0&0\end{bmatrix}^{\top}$, respectively. Notice that this projection corresponds to the Hasse diagram [114].
@@ II.5 Exchanging the labels of vertices
Here, we show a visual understanding of this conversion through contribution basis. By reversing the direction of the contribution basis, one can swap the projected coordinates of the rest, cf. Figs. 2(a) and 2(b), or Figs. 3(a) and 3(b). The dimension of the reversed contribution basis corresponds to the dimension where there is a $1$ in $\bm{b}^{\prime}\oplus\bm{b}^{\prime\prime}$. Visualizing the projected vector provides a way to understand exchanging the labels via reversing the direction of unit vectors.
Figure 3: Exchanging the labels of vertices by reversing the direction of the contribution basis of corresponding digits. The boxes on the bottom right indicate the correspondence between the colors and the dimensions. (a) By reversing the contribution basis for the second digit, one can exchange the cubic labels of pairs of vertices, $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1\end{bmatrix}^{\top}$, and $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Compare this with Fig. 2(a). (b) Obtaining different labels for a tesseract by swapping two contribution basis. The first and fourth unit vectors are reversed, cf. Fig. 2(b).

### v4

@@ Abstract
Projections of hypercubes have been applied to visualize high-dimensional binary state spaces in various scientific fields. Conventional methods for projecting hypercubes, however, face practical difficulties. Manual methods require nontrivial adjustments of the projection basis, while optimization-based algorithms limit the interpretability and reproducibility of the resulting plots. These limitations motivate us to explore theoretically analyzable projection algorithms such as principal component analysis (PCA). Here, we investigate the mathematical properties of PCA-projected hypercubes. Our numerical and analytical results show that PCA effectively captures polarized distributions within the hypercubic state space. This property enables the assessment of the asymptotic distribution of projected vertices and error bounds, which characterize the performance of PCA in the projected space. We demonstrate the application of PCA to visualize the hypercubic energy landscapes of Ising spin systems. By adding projected hypercubic edges, these visualizations reveal pathways of correlated spin flips. We confirm that the time-integrated probability flux exhibits patterns consistent with the pathways identified in the projected hypercubic energy landscapes. Using the mean-field model, we show that dominant state transition pathways tend to emerge around the periphery of the projected hypercubes. Our work provides a better understanding of how PCA discovers hidden patterns in high-dimensional binary data.
@@ I Introduction
By generalizing the idea of a three-dimensional cube to higher dimensions, one obtains a high-dimensional cube or a hypercube [1] (see Appendix A for details on the construction of a hypercube). Although a hypercube is purely a geometric concept [2, 3, 1, 4], it has applications not only in physics [5, 6, 7], electrical engineering [8, 9, 10, 11, 12, 13], graph theory [14, 9], and recreational mathematics [9, 4, 15], but also in a wide range of interdisciplinary arts [16]. Since the pioneering work of almost 150 years ago [2] and subsequent contributions [17, 1], interest in high-dimensional geometry, especially hypercubes, has been continuously growing.
@@
In the sciences, hypercubes are used to visualize binary state111 Also called microstate, configuration, or phase. spaces or state transitions in a wide range of fields, from physics to biology.222 See, for instance, evolutionary landscapes [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], epistasis [18, 30, 31], chemical reaction networks [32, 33], learning in neural networks [34], genetic code space [35, 36, 27], allostery [37, 38, 39, 40, 41, 42], quantum states [43, 44], quantum walks [45], language space [46], data visualization [47, 48], probability currents [49], protein folding [50, 51], energy landscapes [52], gene regulation [53, 54, 55, 56, 57], quantum many-body scar states [58, 59, 60], quantum many-body localization [61, 62, 63, 64, 65, 66, 67, 68], morphology [69, 70], disease progression [27, 28], gene regulatory networks [71], gene expression [72], and neural spike dynamics [73, 74]. This is because hypercubic vertices correspond to binary states of a system [Figs. 1(a) and 1(b)], and hypercubic edges represent transitions333 Strictly speaking, only one of the dimensions, components, or elements of the system must change in the transition on the hypercubic edge. This kind of dynamics is called asynchronous dynamics (update) [75, 76, 77, 71]. It is also called multipartite dynamics [78, 79, 73, 74] in the context of stochastic thermodynamics. On the other hand, transitions with several changes of elements are called synchronous dynamics (update) [80, 81, 82, 83, 84, 85, 86]. between states [Figs. 1(c) and 1(d)]. This can be naturally applied to illustrate the high-dimensional binary state space or state transition diagram. We refer to such a binary state space as a hypercubic state space, and such a state transition diagram as a hypercubic state transition diagram. As an example of transitions between the vertices, we visualize the Hamiltonian path on hypercubes [Figs. 1(c) and 1(d), Tables 1 and 2]. The advantage of visualizing hypercubes is that, in addition to displaying the state space as hypercubic vertices, it also illustrates state transitions as hypercubic edges. Such representations provide valuable insight into the dynamics of many-body systems, particularly by illustrating correlated state transition pathways.
@@
An alternative method to project a hypercube is optimizing the coordinates of projected vertices following a predefined error function [87]. While this method projects hypercubes automatically, it has several limitations. The first limitation is reproducibility. Because of the stochastic nature of the optimization process, projections generated from the same vertices are not guaranteed to be the same (unless all relevant parameters are fixed). Interpretability is another weakness of projecting hypercubes by optimization. Because of its nonlinear nature, what is indicated from the resulting plot is not obvious (see, however, Refs. [88, 89, 90]). How can one create reproducible and interpretable projections of hypercubes?
Here, we reveal that principal component analysis (PCA) [91, 92] provides reproducible, interpretable, and automatic projections of hypercubes. PCA is a linear dimensionality reduction method frequently performed in statistics and machine learning [93, 94]. By interpreting the principal component (PC) loading as the basis for the projection of a hypercube, we show that one can draw the edges of a hypercube on a projected two-dimensional plane. This idea is closely related to the biplot [95], a method to visualize loadings (eigenvectors) with data points to assist in the interpretation of resulting plots. By combining ideas from geometry and statistics (or machine learning), we achieve informative projections of hypercubes [69]: we can analytically obtain the properties of the projection, and the resulting plots are examined with such prior knowledge. Such prior knowledge enhances our interpretation of the resulting plots, enabling a deeper understanding of high-dimensional systems. As an application of these informative projections of hypercubes, we visualize the state space of the Ising spin system and demonstrate that the dynamical behavior of the system can be inferred from the resulting plots.
This paper is organized as follows. In Sec. II, we present the Hamming and fractal projections with introducing manual orthogonal projections of hypercubes. Then, in Sec. III, we show how biplots enable us to interpret the resulting plots of PCA on hypercubic vertices. We provide some examples of orthogonal projections of hypercubes using PCA in Sec. IV. Through analytical and numerical investigations in Sec. V, we show that principal components are informative for weighted vertices. In Sec. VI, using the inner-product error, we show that the vertices around the origin of the projected space less accurately preserve the original distances between them. We apply our method to visualize the hypercubic energy landscapes and probability flux of Ising spin systems in Sec. VII with analytical understanding through the mean-field model, and conclude this paper in Sec. VIII.
@@ II Orthogonal Projections of Hypercubes
Orthogonal projection is a linear method to project a high-dimensional object. We introduce the concept of the contribution vector, which is the projected unit vector of each dimension. We then illustrate isometric, Hamming, and fractal projections of hypercubes by manually varying the contribution basis.
@@ II.1 Dimensionality reduction and orthogonal projection
where $\bm{b}\coloneqq\begin{bmatrix}b_{1}&\cdots&b_{N}\end{bmatrix}^{\top}\in\left\{1,0\right\}^{N}$ is an $N$-dimensional coordinate of the hypercubic vertices. The reduced original dimensions are represented by two $N$-dimensional vectors, $\bm{v}_{1},\bm{v}_{2}\in\mathbb{R}^{N}$. We can write Eq. (1) using the transformation matrix $\bm{V}\coloneqq\begin{bmatrix}\bm{v}_{1}&\bm{v}_{2}\end{bmatrix}\in\mathbb{R}^{N\times 2}$,
$\bm{r}=\bm{V}^{\top}\bm{b}.$ (2)
II.2 Contribution vector
@@
$\bm{r}=\sum_{i=1}^{N}b_{i}\tilde{\bm{e}}_{i},$ (3)
where the contribution vector,
$\tilde{\bm{e}}_{i}\coloneqq\begin{bmatrix}v_{i;1}\\ v_{i;2}\end{bmatrix},$ (4)
@@
[$\ \tilde{\bm{e}}_{1}&\cdots&\tilde{\bm{e}}_{N}\\ \vrule&&\vrule\end{bmatrix}\coloneqq\begin{bmatrix}\text{---}&\bm{v}_{1}^{\top}&\text{---}\\ \text{---}&\bm{v}_{2}^{\top}&\text{---}\end{bmatrix}.$ (5)
@@
The interpretation of the contribution vector in Eq. (3) is the contribution from the $i$th dimension to the projected space. Because $b_{i}\in\left\{1,0\right\}$, the contribution vector corresponds to the projected unit vector of each dimension. Indeed, the alternative definition of contribution vector is
$\tilde{\bm{e}}_{i}\coloneqq\bm{V}^{\top}\bm{e}_{i}$ (6)
where $\bm{e}_{i}\coloneqq\begin{bmatrix}0&\cdots&0&1&0&\cdots&0\end{bmatrix}^{\top}\in\left\{1,0\right\}^{N}$ is the $i$th standard unit vector in the original space. Thus, by introducing the contribution basis $\left\{\bm{e}_{i}\right\}_{i=1}^{N}$, we can decompose the resulting projection into the contributions from each dimension as shown in Eq. (3). The remaining question is how to determine the contribution basis $\left\{\tilde{\bm{e}}_{i}\right\}_{i=1}^{N}$.
Figure 2: A gallery of orthogonal projections of hypercubes. Colored arrows represent the contribution vectors corresponding to the original dimensions. The boxes on the bottom right indicate the correspondence between the colors and the original dimensions. (a) An isometric projection of a cube. Notice that $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$ are overlapped. (b) An isometric projection of a tesseract. (c) A Hamming projection of a cube. (d) A Hamming projection of a tesseract. Notice that $\begin{bmatrix}1&0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1&0\end{bmatrix}^{\top}$ are overlapped. In Hamming projections (c) and (d), the contribution vector of each dimension has the same horizontal contribution. (e) A fractal projection of a six-dimensional hypercube. (f) A fractal projection of an eight-dimensional hypercube. In fractal projections (e) and (f), the contribution vectors of the first half (left half) of the code are ten times longer than the rest. More projections of hypercubes are available in the Supplemental Material [16].
@@ II.4 Hamming and fractal projections
By modifying the contribution basis $\left\{\tilde{\bm{e}}_{i}\right\}_{i=1}^{N}$ for projection, one can view a high-dimensional hypercube from various angles. Several methods reflecting the cross-section of the hypercube have already been suggested [1]. Here, we introduce two other orthogonal projections of hypercubes used in sciences.
A Hamming projection [Figs. 2(c) and 2(d)], we call, is a type of projection that plots vertices according to the Hamming distance [99] from a selected vertex. To achieve this, an example of a contribution vector is $\tilde{\bm{e}}_{i}=\begin{bmatrix}1&\left(i-\frac{N+1}{2}\right)\frac{2}{N-1}\sin\left(\frac{\varphi}{2}\right)\end{bmatrix}^{\top}$, where $\varphi\coloneqq\arccos\left(\frac{\tilde{\bm{e}}_{1}^{\top}\tilde{\bm{e}}_{N}}{\sqrt{\tilde{\bm{e}}_{1}^{\top}\tilde{\bm{e}}_{1}}\sqrt{\tilde{\bm{e}}_{N}^{\top}\tilde{\bm{e}}_{N}}}\right)$ is the angle between the first and last contribution vector, but it can be any. The first element of $\tilde{\bm{e}}_{i}$ is the same for all $i$. The second element can be any, but we determine them to increase linearly with $i$. This type of projection is found not only in research in physics444 Examples are found in quantum walk [45], unsupervised learning of states of Ising spin system [100, 101, 102], quantum many-body scar states [58, 59, 60], and quantum many-body localization [61, 63, 64, 65, 66, 67, 68]. but also in biology.555 Examples include evolutionary landscapes [18, 22, 23, 24, 25, 26, 28, 29], epistasis [18, 30], genetic code space [27], data visualization [47, 48], protein folding [50, 51], gene regulation [53], disease progression [27, 28], and allostery [41]. The horizontal distance between the Hamming projected hypercubic vertices corresponds to the Hamming distance from a reference vertex on the rightmost or leftmost side of the projected space. In Figs. 2(c) and 2(d), the horizontal axis corresponds to the Hamming distance from the origin $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&0&0&0\end{bmatrix}^{\top}$, respectively. Notice that this projection corresponds to the Hasse diagram [103].
@@ II.5 Exchanging the labels of vertices
Here, we show a visual understanding of this conversion through contribution vectors. By reversing the direction of the contribution vector, one can swap the projected coordinates of the rest, cf. Figs. 2(a) and 2(b), or Figs. 3(a) and 3(b). The dimension of the reversed contribution vector corresponds to the dimension where there is a $1$ in $\bm{b}^{\prime}\oplus\bm{b}^{\prime\prime}$. Visualizing the projected vector provides a way to understand exchanging the labels via reversing the direction of unit vectors.
Figure 3: Exchanging the labels of vertices by reversing the direction of the contribution vectors of corresponding digits. The boxes on the bottom right indicate the correspondence between the colors and the dimensions. (a) By reversing the contribution vector for the second digit, one can exchange the cubic labels of pairs of vertices, $\begin{bmatrix}0&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}1&0&0\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&0\end{bmatrix}^{\top}$, $\begin{bmatrix}0&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}0&1&1\end{bmatrix}^{\top}$, and $\begin{bmatrix}1&0&1\end{bmatrix}^{\top}$ and $\begin{bmatrix}1&1&1\end{bmatrix}^{\top}$. Compare this with Fig. 2(a). (b) Obtaining different labels for a tesseract by swapping two contribution vectors. The first and fourth unit vectors are reversed, cf. Fig. 2(b).